OpenStudy (anonymous):

1.Write a function that models the total number of births and deaths in the United States for each year from 2000 through 2005.

OpenStudy (anonymous):

Assume C no of people alive B people being born D people dying then Population P is P = C + B - D. Thus P(2000) = C(2000) + B(2000) - D(2000) which are the counts for that year.